# V2X Cooperative Perception for Autonomous Driving: Recent Advances and Challenges

### v3

@@ V-C Lossy Communication
Drawing inspiration from image-denoising networks, the Lossy Communication Aware Repair Network (LCRN) [228] was designed to generate tensor-wise filters and recover damaged features received from cooperative CAVs. The LCRN employs an encoder-decoder architecture with skip connections to produce per-tensor filter kernels, which are then applied to the received lossy features. The recovered intermediate features subsequently undergo feature fusion through a V2V attention module. The LCRN effectively mitigates the impact of artifacts resulting from damaged features by employing tensor-wise filtering, although it does not explicitly address feature loss. In contrast, the V2X-INCOP [229] leverages historical information to recover missing data caused by interruptions in V2X communication. The core of this approach is an adaptive communication model capable of extracting multi-scale and spatial-temporal features from historical data to estimate missing information. Empirical results have demonstrated that V2X-INCOP can effectively alleviate the adverse effects of communication interruptions on CP.
V-D Data Trustworthiness
Advanced onboard sensors in intelligent vehicles typically provide a high level of reliability for the overall perception system, allowing the vehicle to be aware of its surroundings. However, these sensors are not immune to challenging scenarios like adverse weather conditions and equipment malfunctions. Similarly, various agents within the V2X communication network exhibit changing performances and may possess different equipment grades. Consequently, inconsistencies in transmitted information are inevitable during V2X CP processes. Furthermore, malicious agents may tamper with shared packets or generate forged data [255]. Ego vehicles lack the capability to independently verify events they might have missed and ascertain the authenticity of received data [256]. Overreliance on such misleading information by driver assistance systems can disrupt their functioning, posing risks to road users. Therefore, assessing the quality of received information and detecting erroneous data before proceeding with further processing is essential.
To address this concern, probabilistic modelling techniques were employed in [230] to estimate the trustworthiness of a pair of connected agents based on the consistency of the data they provide. The underlying concept involved quantifying the consistency of received data using the Bayes filter, enabling the identification of forged information based on the relationship between data consistency and agent trustworthiness. However, this method can only detect attackers when there is a substantial redundancy (at least triple redundancy) in the contradictory data.
@@ V-D Data Trustworthiness
Recognizing that spoofing attacks often target a single vehicle at a time, FDII [231] capitalized on LiDAR data from neighbouring vehicles to counteract spoofing attacks on CAVs. By comparing the victim vehicle’s LiDAR scan with unaltered scans from others, FDII could discern the additional points introduced in the victim’s LiDAR data. The differences in detected obstacles under each scan were analyzed using a decision tree algorithm to classify various attacks and update the unsafe areas for each vehicle. The results demonstrated FDII’s effectiveness in accurately identifying attack types and reconstructing vulnerable areas. However, it is worth noting that decision trees may struggle to capture complex data relationships and tend to favour the majority class in imbalanced datasets, potentially leading to suboptimal classification of minority classes.
@@
To mitigate data fabrication attacks, the Collaborative Anomaly Detection (CAD) model [232] enabled benign vehicles to jointly reveal malicious fabrication by sharing occupancy maps which are then merged into one. The occupancy maps are represented using fine-grained polygons, which provide greater flexibility and precision compared to grid-based occupancy maps. Motion estimation is also performed to track objects over consecutive frames and attach motion information to the occupancy map. The consistency checks are then implemented to reveal inconsistencies among synchronized occupancy maps and compare the perception results with the merged occupancy map. Alerts are raised on conflicted regions, indicating potential anomalies or attacks. Although CAD has been shown to be effective in detecting and mitigating attacks in real-world scenarios, it has certain limitations, such as its effectiveness depending on the presence of at least one benign CAV observing the attacked region.
@@
Malicious Agent Detection (MADE) [233] was designed for multi-agent cooperative perception networks, leveraging two statistics, match loss and collaborative reconstruction loss, to detect malicious agents. The match loss measures the difference in output bounding box proposals between the ego agent and the agent being inspected, while the collaborative reconstruction loss assesses the consistency of feature maps when the ego agent collaborates with the agent being inspected. These statistics can be used independently or jointly to construct a multi-test with a controlled false detection rate using the Benjamini-Hochberg procedure [257]. This study provided valuable insights into the detection of malicious agents. However, the detection inference in MADE is based on a multi-test using conformal p-values computed on calibration sets. This approach relies on the assumption that the ego agent is guaranteed to be benign, which may not always hold true in real-world scenarios.
@@ V-E Perception Uncertainty
Multi-agent cooperation significantly enhances the perception efficiency of autonomous vehicles. However, inherent uncertainties persist in CAVs due to out-of-range objects, sensor errors, and adverse weather conditions [258]. These uncertainties can result in false detections, adversely affecting subsequent self-driving modules like planning and control. Addressing these uncertainties within CP is essential for ensuring safe autonomous driving.
@@
Accurately capturing the uncertainties of the perception system represents a promising approach to enhancing CAV safety. The Perception Error Model (PEM) [259] introduced single-agent perception errors into the ground truth to directly model perception uncertainties within the simulation pipeline, bypassing the need for synthetic sensor data. In an expansion of this concept, the Cooperative Perception Error Model (CoPEM) [234] was introduced to tackle occlusion issues by modelling V2X-relevant perception errors and uncertainties. CoPEM facilitates the tuning of perception performance in simulation and scales effectively with an increasing number of sensors for CP. However, it necessitates a statistical model for perception errors for each connected agent and does not account for the high maneuverability of vehicle targets. To address this limitation, the Consensus-based Distributed Multi-Model CP (DMMCP) [235] approach was proposed. DMMCP enhances the model uncertainty of maneuvering targets by integrating prior knowledge from diverse motion models. It employs a hybrid consensus strategy within a multi-layer information fusion framework, combining it with an advanced Cubature Kalman Filter (CKF) [260] to incorporate more precise nonlinear system models and sensor data. DMMCP demonstrates significant effectiveness, particularly when considering model-related uncertainties. Nevertheless, initializing the CKF can be challenging, especially when dealing with uncertainties in the initial state estimate.
While CoPEM and DMMCP provide probabilistic perception results that offer partial information, the GevBEV [171] presents a comprehensive probabilistic model. GevBEV captures perception uncertainty by constructing a probabilistic BEV map of the 2D driving space using point-based spatial Gaussian distributions. It leverages Evidential Deep Learning (EDL) [184] to estimate probabilities and associated uncertainties for new sample points in the driving space. The model interprets the sampled Gaussian distribution densities as evidentiary support for the Dirichlet distribution. Consequently, GevBEV provides a quantifiable and explainable level of uncertainty for its predictive outcomes. Experimental results indicate that GevBEV outperforms previous research on the benchmark OPV2V for BEV map interpretation.
Uncertainty quantification is critical for safety-critical systems in CAVs, helping determine the likelihood of specific outcomes when certain system aspects are imprecisely known. The Double-M quantification method [236] estimates the uncertainty in cooperative object detection by leveraging temporal features in the dataset. At its core, Double-M employs a customized Moving-Block Bootstrap (MBB) algorithm to model the multivariate Gaussian distribution of each corner of the bounding box. MBB accounts for auto-correlations within time-series data by sampling data from constructed data blocks during training. Double-M quantification assigns significant uncertainties to low-accuracy objects, contributing to later modules in self-driving tasks. While Double-M reduces uncertainty and enhances accuracy, it relies on the assumption that each corner of the bounding box adheres to an independent multivariate Gaussian distribution.
In the context of object tracking, the MOT-CUP framework [237] was designed to address uncertainty quantification for cooperative object detection. MOT-CUP incorporates detection uncertainty into the Kalman Filter and the association procedure for object tracking. It quantifies uncertainty in cooperative object detection through direct modelling [261] and conformal prediction [262]. This obtained uncertainty information is then propagated to the Standard Deviation-based Kalman Filter (SDKF) and the Negative Log-likelihood (NLL) based association improvement procedure for multiple object tracking. SDKF uses the uncertainty information to improve location prediction precision and track state prediction. NLL-based metrics are introduced to consider the uncertainty of collaborative detection while handling low-quality detection results. Another 3D multi-object cooperative tracking algorithm, DMSTrack, that utilizes a differentiable multi-sensor Kalman Filter to estimate measurement uncertainty for each detection, was introduced by [238]. The Kalman Filter prediction step used a constant velocity model as the process model. A covariance neural network was designed to learn the observed noise covariance for each detection from a different CAV by using local point cloud features and positional features as inputs. DMSTrack estimates measurement uncertainty for each detection to better utilize the theoretical optimality property of Kalman Filter-based tracking methods. Extensive experiments demonstrate the potential benefits of incorporating uncertainty quantification from CP into multiple object tracking.
@@ V-F Task Discrepancy
V-G Data Privacy
Despite the promising outcomes of collaboration among multiple agents, it necessitates local data exchange among cooperating partners. In real-world scenarios, connected agents often exhibit heterogeneity, and privacy concerns arise as automotive manufacturers and software providers are typically reluctant to share private data. This reluctance results in incomplete and insufficient information for perception tasks. Addressing privacy issues in collaborative driving automation is thus of paramount practical significance.
Conventional FL approaches have limitations in handling data heterogeneity stemming from diverse sensor configurations. Variations in camera positions across different clients can lead to FL results that deviate from the optimal values of local data. To address this challenge, the FedBEVT framework [242] introduces a federated transformer learning approach grounded in camera-attentive personalization. This approach mitigates the adverse effects of data heterogeneity by privatizing the camera positional embedding. Empirical investigations have demonstrated the efficacy of FL techniques in training a transformer-based model to enhance BEV perception. Nevertheless, FedBEVT does not account for network heterogeneity. Varied connection qualities among clients can lead to communication delays during global model aggregation, hindering the FL process.
To tackle the complexities of data and network heterogeneity in vehicular networks, the work in [243] proposed a contextual client selection pipeline. This pipeline leverages digital twin simulations to emulate networks and calculates FL communication latency based on predictive transport scenarios. Specifically, it groups CAVs with similar data distribution and employs the Fast-$\gamma$ rule to select a subset of clients ($\gamma$ clients) with the lowest communication delays per cluster. This FL approach effectively addresses both data and network heterogeneity, enhancing communication efficiency through contextual client selection.
Challenges in FL for CAVs also include unfairness due to differences in data volumes and communication delays. An FL algorithm to address challenges related to imbalanced data distribution and fluctuating channel conditions in CAVs was introduced in the work [241]. A local training strategy was designed specifically for each CAV based on their data volume and channel conditions to reduce the number of global rounds required and shorten the time intervals between each round. This allows it to achieve fairness in energy and time costs, accelerating model convergence while maintaining a balanced cost for all CAVs.
V-H Simulation to Reality
@@ VI-A1 CARLA
CARLA [268] is a pioneering autonomous driving simulator celebrated for its versatility in specifying sensor suites and environmental conditions. This open-source platform was purpose-built to facilitate development, training, and validation of novel autonomous driving systems. CARLA’s architecture is notably scalable, employing a server-multi-client model to distribute computations across multiple nodes. The server continuously updates the environmental physics while users exert control over the client side through the CARLA API. Although CARLA permits environment extension, the process of swiftly creating scenarios and managing substantial traffic volumes has grown intricate. CARLA does offer a traffic manager module for generating background traffic. However, these modules rely on simplified behavioural rules that may not faithfully replicate actual driver behaviour.
@@ VI-A5 CMM Co-simulation Platform
TABLE VIII: Comparison between Cooperative Perception-related Datasets.
Dataset Source Scenario Sensors Tasks
Camera LiDAR Depth IMU/GPS Object Detection Object Tracking Segmentation Motion Prediction Accident Prediction
DAIR-V2X-C [273] Real-world V2I $\surd$ $\surd$ $\times$ $\surd$ $\surd$ $\times$ $\times$ $\times$ $\times$
V2X-Seq [274] Real-world V2I $\surd$ $\surd$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\surd$ $\times$
V2V4Real [275] Real-world V2V $\surd$ $\surd$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\times$
LUCOOP [276] Real-world V2V $\times$ $\surd$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\times$
TUMTrafV2X [277] Real-world V2I $\surd$ $\surd$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\surd$ $\surd$
HoloVIC [278] Real-world V2X $\surd$ $\surd$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\surd$ $\times$
CODD [279] CARLA V2V $\times$ $\surd$ $\times$ $\times$ $\surd$ $\surd$ $\times$ $\times$ $\times$
OPV2V [280] CARLA & OpenCDA V2V $\surd$ $\surd$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\times$
V2X-Sim [281] CARLA & SUMO V2X $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\times$ $\times$
V2XSet [198] CARLA & OpenCDA V2X $\surd$ $\surd$ $\times$ $\times$ $\surd$ $\surd$ $\times$ $\times$ $\times$
DOLPHINS [282] CARLA V2X $\surd$ $\surd$ $\times$ $\times$ $\surd$ $\times$ $\times$ $\times$ $\times$
DeepAccident [283] CARLA V2X $\surd$ $\surd$ $\times$ $\times$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$

### v4

@@ V-B Lossy Communication
Drawing inspiration from image-denoising networks, the Lossy Communication Aware Repair Network (LCRN) [235] utilized an encoder-decoder architecture with skip connections to generate tensor-wise filters for repairing damaged features from lossy communication. These filters are applied to lossy features, which are then fused via a V2V attention module. While LCRN mitigates the impact of feature degradation, it does not directly address feature loss. In contrast, V2X-INCOP [236] recovered missing data by leveraging historical information, using an adaptive communication model that extracts multi-scale spatial-temporal features to estimate missing information, effectively countering communication interruptions. In [243], a Historical Frame Prediction (HFP) module was introduced to address information loss or delay issues by predicting the current frame with temporal associations, taking the stacked results of the previous few historical frames as input.
Previous approaches often overlook realistic channel models and may not capture the complexity of real-world communication channels. To address this, [237] incorporated Rician fading with free-space path loss and the WINNER II model [270] with multi-path fading to simulate practical channels. A self-supervised adaptive weighting model, based on a CNN with a Softmax layer, was introduced to mitigate channel distortion effects on intermediate fusion. The model was trained using contrastive self-supervision without manual annotations, and its generalization was validated across various data domains, detection backbones, noise levels, and path loss factors.
Accessing historical information can reduce transmission frequency and provide a compensatory mechanism for potential communication failures, thereby improving system reliability and performance. Building on this concept, TempCoBEV [238] introduced an importance-guided attention architecture to prioritize critical regions from the historical context, compensating for communication failures. This temporal module can be integrated into state-of-the-art camera-based CP models without requiring complete model re-training, significantly reducing training time. Experiments on the OPV2V dataset demonstrate that TempCoBEV outperforms non-temporal models in predicting current and future BEV map segmentations, particularly in scenarios involving communication failures.
V-C Communication Delays
Latency resulting from network congestion reduces perception accuracy and disrupts fusion alignment [271]. A basic solution to address latency issues is to adjust the timestamp by compensating for the time difference with a reference agent. Building on this concept, CooperFuse [239] introduced a multi-agent temporal synchronization module, incorporating time calibration communication, time compensation, and calibration completion to minimize network transmission latency.
The time series prediction strategy is also effective in addressing the latency issue by predicting the sender’s perception at the current time step based on historical frames. In this context, the FFNet [240] leveraged temporal coherence by extracting the first-order derivative of feature flow from previous infrastructure frames to predict future features and mitigate asynchrony. While effective for very short time intervals, this first-order expansion method has limitations. SyncNet [241] adopted a dual-branch pyramid Long Short-term Memory (LSTM) network to infer real-time features and cooperative attention. The HFP module [243] can generate the prediction of the current frame through a series of multidimensional convolution mixing operations. However, these approaches have not effectively addressed issues caused by irregular time delays. As a substitute, CoBEVFlow [242] treated each collaboration message as an irregular sample, using historical frames to estimate the BEV flow map for asynchronous feature alignment, though it still depends on past data for latency compensation. In contrast, V2X-PC [136] bypassed this reliance by directly utilizing low-level coordinate information in point clusters to predict their locations at the current timestamp.
Re-encoding received features with delay time is another approach for mitigating communication delays. For instance, V2X-ViT [145] addressed time delay through a Delay-aware Positional Encoding module that utilizes sinusoid functions conditioned on delay time and channel information to initialize a learnable embedding. Similarly, V2VNet [140] utilized a CNN to manage latency by incorporating the received intermediate feature, relative 6DoF pose, and delay time, producing a time-delay-compensated representation. These methods assume latency is known; however, latency is typically calculated using timestamps from different agents’ clocks, which may result in deviations and inaccurate latency measurements. To address this, FreeAlign [195] introduced a multi-anchor based subgraph searching algorithm to identify the common subgraph between salient-object graphs across different agents, estimating clock deviations and determining the precise time difference between collaborative messages. This approach ensures accurate temporal alignment without relying on external clock signals.
V-D Security and Privacy
Data security and privacy are essential to the reliability of CP systems. V2X communication networks connecting various agents are vulnerable to potential attacks during data transmission, which could compromise system integrity. Furthermore, data sharing between agents introduces risks of privacy breaches, underscoring the need for robust security protocols within CP networks.
V-D1 Data Security
While cooperative multi-agent systems offer significant promise, they also present security risks due to the potential for malicious or unreliable communication between agents [272]. Malicious agents may tamper with shared packets or generate forged data. The ego vehicle is unable to independently verify missed events or confirm the authenticity of received data [273]. Imperceptible perturbations introduced by adversaries can substantially alter perception outputs, compromising system integrity. Although adversarial training has been used to address these threats [244], it adds training overhead and often fails to generalize to novel attacks [274]. As a solution, ROBOSAC [245] was proposed, allowing ego vehicles to intelligently select trustworthy collaborators. By detecting divergences in adversarial messages and aligning with reliable collaborators, ROBOSAC generalizes to unseen adversarial strategies and offers a flexible trade-off between performance and computational cost.
To ensure data security, assessing the quality of received information and detecting erroneous data before processing is crucial. Probabilistic modelling techniques have been utilized to estimate the trustworthiness of connected agents based on data consistency [246]. This approach applies a Bayes filter to quantify consistency, allowing the detection of forged information by analyzing the correlation between data consistency and agent reliability. However, this method is constrained in detecting attackers, as it requires substantial redundancy—specifically, at least triple redundancy in conflicting data—to effectively identify malicious agents.
@@ V-D1 Data Security
FDII [247] addressed spoofing attacks by leveraging LiDAR data from neighbouring vehicles to detect tampered points in the victim’s scan. By comparing the altered and unaltered scans, FDII identifies discrepancies and classifies attacks using a decision tree algorithm, which also updates the unsafe areas for each vehicle. The approach proved effective in identifying attack types and reconstructing vulnerable regions. However, decision trees may struggle with complex data relationships and often a bias toward the majority class in imbalanced datasets, potentially leading to the suboptimal classification of minority classes.
@@
To counter data fabrication attacks, the Collaborative Anomaly Detection (CAD) model [248] allows benign vehicles to collectively detect malicious fabrications by sharing and merging occupancy maps. These maps, represented by fine-grained polygons, offer greater precision and flexibility compared to grid-based alternatives. CAD incorporates motion estimation to track objects across frames, attaching motion data to the maps. Consistency checks are performed to identify discrepancies among synchronized maps and compare perception results with the merged map, raising alerts for conflicted regions that suggest potential anomalies. Although CAD proves effective in real-world scenarios, its efficacy depends on the presence of at least one benign CAV monitoring the attacked region.
@@
Malicious Agent Detection (MADE) [249] used two key statistics, match loss and collaborative reconstruction loss, to identify malicious agents. Match loss quantifies discrepancies in bounding box proposals between the ego agent and the inspected agent, while collaborative reconstruction loss measures the consistency of feature maps when the two agents collaborate. These metrics can be applied independently or together to form a multi-test with a controlled false detection rate, utilizing the Benjamini-Hochberg procedure [275]. Although this approach offers valuable insights, MADE’s reliance on conformal p-values computed from calibration sets assumes the ego agent is benign, a condition that may not always be valid in real-world scenarios.
V-D2 Data Privacy
While collaboration among multiple agents shows great potential, it requires local data exchange between cooperating partners. In practice, connected agents often demonstrate heterogeneity and privacy concerns emerge as automotive manufacturers and software providers are generally unwilling to share proprietary data. This reluctance leads to incomplete and insufficient information for perception tasks, making it crucial to address privacy issues in V2X CP systems.
Conventional FL approaches face challenges in managing data heterogeneity arising from diverse sensor configurations. Variations in camera positions across clients can result in FL results that deviate from the optimal values of local data. To tackle this issue, the FedBEVT [251] framework introduced a federated transformer learning approach focused on camera-attentive personalization, which alleviates the negative impacts of data heterogeneity by privatizing the camera positional embedding. The work in [253] addressed the data heterogeneity in CP by implementing a novel federated dynamic weighted aggregation (FedDWA) algorithm along with a dynamic adjusting loss function. However, these methods ignore network heterogeneity. Varying connection qualities among clients can result in communication delays, thereby impeding the FL process. To fill this gap, a contextual client selection pipeline was introduced in [252], modelling networks and estimating FL communication latency under predictive transport scenarios via digital twin simulations. This FL methodology effectively tackles data and network heterogeneity, thereby improving communication efficiency through contextual client selection.
Challenges in FL for CAVs also include unfairness due to differences in data volumes and communication delays. To address the issues of imbalanced data distribution and variable channel conditions, the study in [250] employed a local training strategy customized for each CAV based on its data volume and channel conditions, thereby minimizing the number of global rounds needed and reducing the time intervals between rounds. Consequently, it promotes fairness in energy and time expenditure, accelerating model convergence while ensuring equitable costs for all CAVs.
@@ V-E Perception Uncertainty
Multi-agent cooperation markedly enhances the perception efficiency of autonomous vehicles. However, inherent uncertainties persist in CAVs due to out-of-range objects, sensor errors, and adverse weather conditions [279]. These uncertainties can result in false detections, adversely affecting subsequent self-driving modules like planning and control. Addressing these uncertainties within CP is essential for ensuring safe autonomous driving.
@@
Capturing perception system uncertainties is crucial for enhancing CAV safety. The Perception Error Model (PEM) [280] incorporated single-agent perception errors into the ground truth to directly model uncertainties in the simulation pipeline, eliminating the need for synthetic sensor data. Expanding on this, the Cooperative Perception Error Model (CoPEM) [254] addresses occlusion by modelling V2X-related perception errors and uncertainties, facilitating performance tuning as sensor numbers increase. However, CoPEM requires a statistical model for each agent and overlooks the high maneuverability of vehicle targets. To resolve this, the Consensus-based Distributed Multi-Model CP (DMMCP) [255] integrated prior knowledge from diverse motion models and applied a hybrid consensus strategy with a Cubature Kalman Filter (CKF) [281] for more accurate data fusion and model uncertainty handling. Although effective, CKF initialization remains challenging, especially with uncertain initial state estimates. CMP [256] used the AB3DMOT tracker [282] to associate the detected 3D bounding boxes into trajectory segments. By incorporating cooperative perception and motion prediction aggregation, CMP is able to address the issue of propagated uncertainties and inaccuracies from the upstream detection and tracking tasks, leading to improved overall motion prediction performance.
Uncertainty quantification is essential for ensuring safety in CAV systems by estimating the likelihood of outcomes when certain system aspects are uncertain. The Double-M [257] addressed uncertainty in cooperative object detection by utilizing temporal features. It applies a customized Moving-Block Bootstrap algorithm to model the multivariate Gaussian distribution for each bounding box corner, accounting for auto-correlations in time-series data. By assigning higher uncertainties to low-accuracy objects, Double-M improves accuracy, though it assumes each bounding box corner follows an independent multivariate Gaussian distribution. GevBEV [133] introduced a more comprehensive model by constructing a probabilistic BEV map using point-based spatial Gaussian distributions. It leverages Evidential Deep Learning [159] to estimate probabilities and uncertainties, interpreting Gaussian densities as support for a Dirichlet distribution. This allows GevBEV to quantify and explain prediction uncertainty, with experimental results showing it outperforms previous models on the OPV2V benchmark.
In the context of object tracking, the MOT-CUP framework [258] addressed uncertainty quantification by integrating detection uncertainty into the Kalman Filter (KF) and the object association process. It models uncertainty using direct modelling [283] and conformal prediction [284], which is then applied to the Standard Deviation-based KF (SDKF) and a Negative Log-Likelihood (NLL) association improvement procedure. SDKF leverages uncertainty to enhance location prediction accuracy, while NLL improves tracking by accounting for low-quality detections. Similarly, DMSTrack [259] used a differentiable multi-sensor KF to estimate uncertainty for each detection, optimizing tracking through a learned covariance model. Nevertheless, neither approach uses multi-vehicle tracking information for improved data association, and temporary tracking failures remain a challenge. To this end, the work in [260] proposed a complementary data association module to identify and compensate for missed objects using information shared among the CAVs. Experiments show that incorporating uncertainty improves multi-object tracking performance.
@@ V-F Task Discrepancy
V-G Simulation to Reality
@@ VI-A1 CARLA
CARLA [287] is a pioneering autonomous driving simulator celebrated for its versatility in specifying sensor suites and environmental conditions. This open-source platform was purpose-built to facilitate the development, training, and validation of novel autonomous driving systems. CARLA’s architecture is notably scalable, employing a server-multi-client model to distribute computations across multiple nodes. The server continuously updates the environmental physics while users exert control over the client side through the CARLA API. Although CARLA permits environment extension, the process of swiftly creating scenarios and managing substantial traffic volumes has grown intricate. CARLA does offer a traffic manager module for generating background traffic. However, these modules rely on simplified behavioural rules that may not faithfully replicate actual driver behaviour.
@@ VI-A5 CMM Co-simulation Platform
TABLE VII: Comparison between Cooperative Perception-related Datasets.
Category Dataset Year Scenario Agent Number Annotation Range Object Category RGB Image LiDAR Frame 3D Box Maps OD OT SS MP AP
Real-world DAIR-V2X-C [292] 2022 V2I 2 280m 10 39k 39k 464k $\times$ $\surd$ $\times$ $\times$ $\times$ $\times$
V2X-Seq [293] 2023 V2I 2 280m 10 15k 15k -- $\surd$ $\surd$ $\surd$ $\times$ $\surd$ $\times$
HoloVIC [294] 2024 V2I 2 70m 3 100k 100k 11.5M $\times$ $\surd$ $\surd$ $\times$ $\surd$ $\times$
V2V4Real [295] 2023 V2V 2 200m 5 40k 20k 240k $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\times$
LUCOOP [296] 2023 V2V 3 200m 5 0 54k -- $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\times$
TUMTrafV2X [297] 2024 V2I 2 200m 8 5k 2k 30k $\surd$ $\surd$ $\surd$ $\times$ $\surd$ $\surd$
RCooper [298] 2024 I2I 2 230m 10 50k 30k -- $\times$ $\surd$ $\surd$ $\times$ $\times$ $\times$
InScope [299] 2024 I2I 2 230m 4 0 21k 188k $\times$ $\surd$ $\surd$ $\times$ $\times$ $\times$
V2X-Real [300] 2024 V2X 4 200m 10 171k 33k 1.2M $\surd$ $\surd$ $\times$ $\times$ $\times$ $\times$
Similation CODD [301] 2022 V2V 4–16 100m 2 0 13.5k 204k $\times$ $\surd$ $\surd$ $\times$ $\times$ $\times$
OPV2V [302] 2022 V2V 2–7 120m 1 44k 11k 233k $\surd$ $\surd$ $\times$ $\surd$ $\times$ $\times$
V2X-Sim [303] 2022 V2X 2–5 70m 1 60k 10k 26k $\surd$ $\surd$ $\surd$ $\surd$ $\times$ $\times$
V2XSet [145] 2022 V2X 2–7 120m 1 44k 11k 233k $\surd$ $\surd$ $\times$ $\times$ $\times$ $\times$
DOLPHINS [304] 2022 V2X 3 200m 2 42k 42k 292k $\times$ $\surd$ $\times$ $\times$ $\times$ $\times$
DeepAccident [305] 2023 V2X 5 70m 6 57k 57k 285k $\times$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
SCOPE [306] 2024 V2X 3–24 250m 5 17.6k 17.6k 575k $\times$ $\surd$ $\times$ $\surd$ $\times$ $\times$
  • Note: OD: Object Detection, OT: Object Tracking, SS: Semantic Segmentation, AP: Accident Prediction.